OpenStudy (anonymous):

How do I find A1 of a geometric sequence given A4=125 and A7=1953.125?

OpenStudy (tkhunny):

\(a_{4} = a_{1}r^{3} = 125\) \(a_{6} = a_{1}r^{6} = 1953.125\)

OpenStudy (anonymous):

So would I then subtract?

OpenStudy (tkhunny):

You would use the two equations in any way you deemed necessary to solve the problem.

OpenStudy (anonymous):

raise both to their respective power, no?

OpenStudy (tkhunny):

You have two equations in two unknowns. \(a_{1}r^{3} = 125\) \(a_{1}r^{6} = 1953.125\) How do you find \(a_{1}\;and\;r\)? It is algebra. You've been preparing for this moment your entire mathematics career. Breakthrough the darkness and see it! How do you find the two values?

OpenStudy (anonymous):

Thanks

OpenStudy (tkhunny):

Why would I be saying anything else? Except, of course, that I never would use the term "plug in". :-)

OpenStudy (anonymous):

substitute?

OpenStudy (tkhunny):

Personally, I might do this: \(a_{1}r^{6} = a_{1}r^{3}r^{3} = 125r^{3} = 1953.125\) and you are a moment away from \(r\).

OpenStudy (anonymous):

I like that, thanks to both of you

OpenStudy (anonymous):

errr.... I have 125r^3=1953.125 I raise both to ^(1/3) and somehow got 125r=12.5??? Where's the error?

OpenStudy (tkhunny):

I was really torn on this problem. We want \(a_{1}\), so it is against my inclination to solve first for \(r\). However, I did want to show you that slightly odd substitution. Keep your minds open. I would prefer to solve directly for \(a_{1}\) I might do it like this: \(r^{6} = \left(r^{3}\right)^{2} = \dfrac{1953.125}{a_{1}} = \left(\dfrac{125}{a_{1}}\right)^{2} \implies a_{1} = \dfrac{125^{2}}{1953.125} = 8\) ...and I NEVER have to know what \(r\) is!!

OpenStudy (tkhunny):

If you must... \(125r^{3}=1953.125\implies r^{3} = \dfrac{1953.125}{125} = 15.625 \implies r = 2.5\)

OpenStudy (anonymous):

r makes sense, let me run through a1 quickly

OpenStudy (anonymous):

awesome, thanks tk!

OpenStudy (tkhunny):

No worries. It's what we do, here. Keep your mind open!! Don't get stuck on one way of thinking!